OpenStudy (anonymous):

the isotopes of carbon (carbon-12,carbon-13,and carbon-14) have difernt retricenumbers each isotopes has six protons. explain how you can determine the number of neutrons in each isotope.

OpenStudy (anonymous):

Im gonna be completely honest with you...just look this question up on the internet and it will come more faster:/ it takes forver here*.* to get ur question anwsered

OpenStudy (anonymous):

Atomic weight = #protons + # neutrons in nucleus atomic number = 6 thus #neutrons = 12-6=6, 13-6=7, 14-6=8